# An Interdisciplinary and Cross-Task Review on Missing Data Imputation

### v2

@@ 5.2 Regression Imputation
The aforementioned zero, mean, median, and mode imputations only consider the distribution of the values of the variable that has missing entries. However, if there is a correlation between the target variable and other variables, which is often the case in practice, regression of the missing variable on these other variables often provides more accurate estimations, especially when data are MNAR (little1988missing; roth1994missing). For instance, in a bivariate analysis where one variable has missing data, the regression equation is estimated using only the cases with complete data. In this equation, the variable with incomplete data is treated as the outcome, while the variable with complete data serves as the predictor. baraldi2010introduction provided an intuitive example that used the math score to predict course grade. More generally, suppose we have $d$ variables $x_{1},x_{2},\ldots,x_{d}$ and we want to impute the missing values for $x_{j}$. Regression imputation first uses the complete observations to build a regression model, such as
@@
An important variation of regression imputation is the iterative regression method (roth1994missing). It begins by calculating an initial correlation matrix using a straightforward imputation method, like mean imputation, or simply using the observed values. Following this, regression equations are derived from this matrix, and the missing values are estimated and replaced in the data matrix. This cycle continues until the regression coefficients exhibit minimal changes.
Perhaps the best-known iterative regression imputation algorithm is the missForest proposed by MissForest. In missForest, the $f^{(j)}$ in (1) is a random forest, and there are $d$ random forests in total. The main steps of missForest are summarized as follows:
@@
The random forest within the missForest algorithm can be replaced with other regression models, such as polynomial regression, support vector regression (drucker1996support), Gaussian processes (williams1995gaussian), or neural networks (lecun2015deep). This substitution can create new regression-based imputation algorithms, though potentially at the cost of higher computational complexity and reduced effectiveness with categorical data. In practice, missForest remains a stable and effective benchmark. To the best of our knowledge, no variant proposed in the past decade has substantially surpassed its imputation accuracy. For instance, while liu2023improving incorporated a Markov Blanket discovery method to refine the feature selection process in missForest, it achieved only a marginal improvement in accuracy.
Regression imputation has several potential drawbacks. It can artificially reduce the variance of imputed variables and introduce bias if the underlying model assumptions are violated (roth1994missing; baraldi2010introduction). Furthermore, if the proportion of missing data is very high (e.g., 90%), estimating the regression function $f^{(j)}$ may be infeasible or highly inaccurate, leading to imputation failure. Although iterative regression methods can mitigate this issue, they remain sensitive to the initial estimates, and an unreliable starting correlation matrix can still yield poor final results.
@@ 5.3 Hot-Deck Imputation
Hot-deck imputation (andridge2010review) replaces the missing values of one or several variables for a non-respondent (known as the recipient) with observed values from a respondent (referred to as the donor) who shares similar characteristics with the non-respondent. Some variations of this method involve randomly selecting the donor from a group of potential donors, also known as the donor pool. These are termed as random hot deck methods. Conversely, some versions identify a single donor, typically the nearest neighbor (NN) based on a certain metric, and impute values from that particular case. These are referred to as deterministic hot-deck methods, as they don’t involve any randomness in the selection of the donor. There are also methods that impute summaries of values from a group of donors, like the mean or weighted average, instead of individual values. The process of hot-deck imputation typically involves the following steps:
@@
This method generally offers greater accuracy than basic imputation strategies, as it replaces missing data with plausible values derived from similar, observed cases rather than a single statistic like the mean, which can distort the underlying variable distributions. A key advantage is that the imputed values preserve the original data’s distributional characteristics because they are conditioned on the matching (categorization) variables. This approach is particularly beneficial when data is missing in specific patterns (roth1994missing). The methodology, implementation, and applications of this technique, known as hot-deck imputation, are detailed in (andridge2010review).
For hot-deck imputation, the nearest neighbors or the donor pool are found by using a distance or similarity metric. There are many choices for the distance or similarity metric. For numerical data, one may use the Euclidean distance, $\ell_{1}$ distance, cosine similarity, Mahalanobis distance, Chebyshev distance, etc (FKIH20227645). For categorical data, one may use the overlap measure (or Hamming distance), Eskin (eskin2002geometric), Goodall’s measures (goodall1966new), Gambaryan measure, inverse occurrence frequency, etc. A review of similarity measures for categorical data can be found in (chandola2007similarity). Consequently, for mixed-type data, one needs to combine different distance or similarity measures.
The most representative example of hot-deck imputation is kNN imputation (chen2000nearest; beretta2016nearest), which is a deterministic hot-deck method. Suppose ${x}_{ij}$, the variable $j$ at observation $i$, is a missing value. kNN imputation make prediction for ${x}_{ij}$ by the following weighted sum
@@
where $\mathcal{N}_{k}$ denotes the index set of the $k$ nearest neighbors of observation $i$ without missing values of variable $j$. $w_{v}$ are the weights of the neighbors and satisfy $\sum_{v\in\mathcal{N}_{k}}w_{v}=1$. If $w_{v}$, $v\in\mathcal{N}_{k}$, are all $1/k$, $\hat{x}_{ij}$ is just the average of the $k$ nearest neighbors. In recommendation systems, $w_{v}$ are usually set to $w_{v}=s_{vi}/\sum_{l\in\mathcal{N}_{k}}s_{li}$ (breeseCF1998; sarwar2001item), where $s_{li}$ denotes the similarity between observations $l$ and $i$. Note that $s_{li}$ is computed from only the overlapped observed values of the two samples.
@@ 5.4 Likelihood Methods
Maximum likelihood is a strategy for estimating the unknown parameters of a model from data. When data is complete, it’s straightforward to calculate quantities like the mean, covariance, and linear regression coefficients, which are maximum likelihood estimates derived from maximizing the likelihood of the observed data. This same principle applies even when there are missing values in the data, as the estimation can still be based on the likelihood of the observed values. However, when missing data is present, the likelihood of the observed data becomes more complex compared to the typical data analysis scenarios, and finding an estimate that maximizes this likelihood doesn’t have a straightforward solution. To tackle the challenge, dempster1977maximum proposed the Expectation-Maximization (EM) algorithm for incomplete data. The main idea and algorithm are explained in the following context.
@@
The two steps are repeated until some convergence conditions are met. Finally, the missing values are filled by using the estimated $\theta$. See the example of multivariate Gaussian in Section 11.6 of (murphy2012machine).
@@ 5.5.1 Low-Rank Matrix Completion
Low-rank matrix completion (LRMC) (srebro2005rank; CandesRecht2009; wen2012solving; hardt2014understanding) addresses the problem of recovering missing entries in a matrix under the assumption that the true matrix is low-rank. Formally, a matrix is low-rank if its rank (the dimension of the vector space spanned by its rows or columns) is significantly smaller than its dimensions. An equivalent characterization is that the matrix has a rapidly decaying singular value spectrum, with only a few non-zero or significant singular values. Consider a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, where $m<n$, the singular value decomposition (SVD) is denoted as $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthonormal matrices, i.e., $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{m}$ and $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{n}$, $\bm{\Sigma}=(\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m}),\bm{0})$, and $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$ are singular values. The rank of $\mathbf{X}$ is defined as $\text{rank}(\mathbf{X}):=\sum_{i=1}^{m}\mathbbm{1}(\sigma_{i}\neq 0)$. It has been found that many data matrices in real applications are low-rank ($\text{rank}(\mathbf{X})\ll m$) or can be well approximated by a low-rank matrix, i.e., many singular values are close to zero.
@@
nuclear norm $\|\mathbf{X}\|_{\ast}=\sum_{i=1}^{m}\sigma_{i}(\mathbf{X})$ (CandesRecht2009)
truncated nuclear norm $\|\mathbf{X}\|_{\text{TN}}=\sum_{i=r+1}^{m}\sigma_{i}(\mathbf{X})$ (hu2012fast)
weighted nuclear norm $\|\mathbf{X}\|_{\text{WN}}=\sum_{i=1}^{m}w_{i}\sigma_{i}(\mathbf{X})$ (gu2014weighted)
Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{Sp}}=(\sum_{i=1}^{m}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (MC_ShattenP_AAAI125165)
weighted Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{WSp}}=(\sum_{i=1}^{m}w_{i}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (7539605)
generalized non-convex penalty $R(\mathbf{X})=\sum_{i=1}^{m}g_{\lambda}(\sigma_{i}(\mathbf{X}))$ (lu2014generalized)
@@
factored nuclear norm $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{F}^{2}+\|\mathbf{B}\|_{F}^{2})$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\ast}$ (srebro2005maximum)
max norm $R(\mathbf{A},\mathbf{B})=(\max_{i}\left\|\mathbf{A}_{i:}\right\|_{2})(\max_{j}\left\|\mathbf{B}_{:j}\right\|_{2})$ srebro2005rank
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{*}+\|\mathbf{B}\|_{*})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (shang2016tractable)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{2,1}+\left\|\mathbf{B}^{T}\right\|_{2,1})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (fan2019factor)
factored Schatten-$2q/(2+q)$ $R(\mathbf{A},\mathbf{B})=\frac{1}{q}\|\mathbf{A}\|_{2,q}^{q}+\frac{\alpha}{2}\left\|\mathbf{B}^{T}\right\|_{F}^{2}$ $\min R(\mathbf{A},\mathbf{B})=c\|\mathbf{X}\|_{S_{2q/(2+q)}}^{2q/(2+q)}$ (fan2019factor)
factored Schatten-$p$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2^{p}}\sum_{i=1}^{d}\left(\left\|\mathbf{A}_{:i}\right\|_{2}^{2}+\left\|\mathbf{B}_{i:}\right\|_{2}^{2}\right)^{p}$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{S_{p}}^{p}$ (giampouras2020novel)
@@
Beyond general low-rank matrices, several LRMC methods are designed for specific types. Notable examples include distance matrix completion (alfakih1999solving), covariance matrix completion (hippert2022robust), kernel matrix completion (tsuda2003algorithm), and one-bit matrix completion (davenport20141). In the latter case, the observed data are binary (1-bit) measurements rather than precise numerical values. This approach is valuable in scenarios with highly quantized data or binary feedback, such as in recommender systems and social networks.
@@ 5.5.2 High-Rank Matrix Completion
where $f^{(1)},\ldots,f^{(s)}$ are different nonlinear functions. This forms a matrix $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\ldots,\mathbf{X}^{(s)}]$ of size $m\times n$, where $\mathbf{X}^{(j)}=[\mathbf{x}_{1}^{(j)},\ldots,\mathbf{x}^{(j)}_{n_{j}}]\in\mathbb{R}^{m\times n_{j}}$ and $n=\sum_{j=1}^{s}n_{j}$. It can be seen that (14) degrades to (13) if $s=1$ and degrades to (12) if all $f^{(j)}$ are linear. Figure 8 provides synthetic examples for (12),(13), and (14), respectively. The completion of the matrix produced by (14) poses the greatest challenge.
Figure 8: Examples (fan2020polynomial) of data forming high-rank matrices in 3D space (left: union of subspaces; middle: one nonlinear manifold; right: union of nonlinear manifolds).
In recent years, a few high-rank matrix completion (HRMC) methods have been proposed (ErikssonBalzanoNowak2011; NLMC2016; NIPS2016_6357; fan2018matrix; fan2020polynomial; NEURIPS2020_42ae1544). They can be organized into three categories: subspace-based methods, kernel-based methods, and deep learning-based methods.
@@
These methods were proposed for the data given by (12) and usually learn the subspaces from the incomplete data (ErikssonBalzanoNowak2011; 8469061) or exploit the self-expressive model (yang2015sparse; NIPS2016_6357; fan2017sparse; FAN2017290), where each data point can be well represented by other data points drawn from the same subspace. The self-expressive model (elhamifar2013sparse) can be formulated as
@@
where $\mathbf{S}\in\mathbb{R}^{n\times n}$ is the coefficient matrix and is usually assumed to be sparse. Given an incomplete data matrix $\tilde{\mathbf{X}}$, FAN2017290 proposed the following matrix completion method:
@@
These methods (NLMC2016; pmlr-v70-ongie17a; FANNLMC; fan2019online; fan2020polynomial; ongie2021tensor) take advantage of kernel techniques that can convert a nonlinear problem in the original data space to a linear problem in the high-dimensional feature space. pmlr-v70-ongie17a and FANNLMC proposed to solve
@@
fan2019online showed that $\text{rank}(\phi(\mathbf{X}))=\min\{\binom{m+q}{q},n,s\binom{d+pq}{pq}\}$ if $f^{(j)}$ are $p$-order polynomials and $\phi$ is the feature map of a $q$-order polynomial kernel. When $n$ is sufficiently large and $p$ and $q$ are relatively small, the rank of $\phi(\mathbf{X})$ is low compared to its dimensions. fan2019online introduced a kernel factorization method for HRMC and an online algorithm, which are more efficient than the methods proposed in (pmlr-v70-ongie17a; FANNLMC).
@@
These methods (FAN2017540; fan2018matrix; seyedi2023elastic) are based on neural networks and will be detailed in Section 5.6.

### v3

@@ 5.2 Regression Imputation
The aforementioned zero, mean, median, and mode imputations only consider the distribution of the values of the variable that has missing entries. However, if there is a correlation between the target variable and other variables, which is often the case in practice, regression of the missing variable on these other variables often provides more accurate estimations, especially when data are MNAR (Little, 1988; Roth, 1994). For instance, in a bivariate analysis where one variable has missing data, the regression equation is estimated using only the cases with complete data. In this equation, the variable with incomplete data is treated as the outcome, while the variable with complete data serves as the predictor. Baraldi and Enders (2010) provided an intuitive example that used the math score to predict course grade. More generally, suppose we have $d$ variables $x_{1},x_{2},\ldots,x_{d}$ and we want to impute the missing values for $x_{j}$. Regression imputation first uses the complete observations to build a regression model, such as
@@
An important variation of regression imputation is the iterative regression method (Roth, 1994). It begins by calculating an initial correlation matrix using a straightforward imputation method, like mean imputation, or simply using the observed values. Following this, regression equations are derived from this matrix, and the missing values are estimated and replaced in the data matrix. This cycle continues until the regression coefficients exhibit minimal changes.
Perhaps the best-known iterative regression imputation algorithm is the missForest proposed by Stekhoven and Bühlmann (2011). In missForest, the $f^{(j)}$ in (1) is a random forest, and there are $d$ random forests in total. The main steps of missForest are summarized as follows:
@@
The random forest within the missForest algorithm can be replaced with other regression models, such as polynomial regression, support vector regression (Drucker et al., 1996), Gaussian processes (Williams and Rasmussen, 1995), or neural networks (LeCun et al., 2015). This substitution can create new regression-based imputation algorithms, though potentially at the cost of higher computational complexity and reduced effectiveness with categorical data. In practice, missForest remains a stable and effective benchmark. To the best of our knowledge, no variant proposed in the past decade has substantially surpassed its imputation accuracy. For instance, while Liu and Constantinou (2023) incorporated a Markov Blanket discovery method to refine the feature selection process in missForest, it achieved only a marginal improvement in accuracy.
Regression imputation has several potential drawbacks. It can artificially reduce the variance of imputed variables and introduce bias if the underlying model assumptions are violated (Roth, 1994; Baraldi and Enders, 2010). Furthermore, if the proportion of missing data is very high (e.g., 90%), estimating the regression function $f^{(j)}$ may be infeasible or highly inaccurate, leading to imputation failure. Although iterative regression methods can mitigate this issue, they remain sensitive to the initial estimates, and an unreliable starting correlation matrix can still yield poor final results.
@@ 5.3 Hot-Deck Imputation
Hot-deck imputation (Andridge and Little, 2010) replaces the missing values of one or several variables for a non-respondent (known as the recipient) with observed values from a respondent (referred to as the donor) who shares similar characteristics with the non-respondent. Some variations of this method involve randomly selecting the donor from a group of potential donors, also known as the donor pool. These are termed as random hot deck methods. Conversely, some versions identify a single donor, typically the nearest neighbor (NN) based on a certain metric, and impute values from that particular case. These are referred to as deterministic hot-deck methods, as they don’t involve any randomness in the selection of the donor. There are also methods that impute summaries of values from a group of donors, like the mean or weighted average, instead of individual values. The process of hot-deck imputation typically involves the following steps:
@@
This method generally offers greater accuracy than basic imputation strategies, as it replaces missing data with plausible values derived from similar, observed cases rather than a single statistic like the mean, which can distort the underlying variable distributions. A key advantage is that the imputed values preserve the original data’s distributional characteristics because they are conditioned on the matching (categorization) variables. This approach is particularly beneficial when data is missing in specific patterns (Roth, 1994). The methodology, implementation, and applications of this technique, known as hot-deck imputation, are detailed in (Andridge and Little, 2010).
For hot-deck imputation, the nearest neighbors or the donor pool are found by using a distance or similarity metric. There are many choices for the distance or similarity metric. For numerical data, one may use the Euclidean distance, $\ell_{1}$ distance, cosine similarity, Mahalanobis distance, Chebyshev distance, etc (Fkih, 2022). For categorical data, one may use the overlap measure (or Hamming distance), Eskin (Eskin et al., 2002), Goodall’s measures (Goodall, 1966), Gambaryan measure, inverse occurrence frequency, etc. A review of similarity measures for categorical data can be found in (Chandola et al., 2007). Consequently, for mixed-type data, one needs to combine different distance or similarity measures.
The most representative example of hot-deck imputation is kNN imputation (Chen and Shao, 2000; Beretta and Santaniello, 2016), which is a deterministic hot-deck method. Suppose ${x}_{ij}$, the variable $j$ at observation $i$, is a missing value. kNN imputation make prediction for ${x}_{ij}$ by the following weighted sum
@@
where $\mathcal{N}_{k}$ denotes the index set of the $k$ nearest neighbors of observation $i$ without missing values of variable $j$. $w_{v}$ are the weights of the neighbors and satisfy $\sum_{v\in\mathcal{N}_{k}}w_{v}=1$. If $w_{v}$, $v\in\mathcal{N}_{k}$, are all $1/k$, $\hat{x}_{ij}$ is just the average of the $k$ nearest neighbors. In recommendation systems, $w_{v}$ are usually set to $w_{v}=s_{vi}/\sum_{l\in\mathcal{N}_{k}}s_{li}$ (Breese et al., 1998; Sarwar et al., 2001), where $s_{li}$ denotes the similarity between observations $l$ and $i$. Note that $s_{li}$ is computed from only the overlapped observed values of the two samples. An illustrative example is shown in Figure 9.
Figure 9: A simple example of kNN imputation.
@@ 5.4 Likelihood Methods
Maximum likelihood is a strategy for estimating the unknown parameters of a model from data. When data is complete, it’s straightforward to calculate quantities like the mean, covariance, and linear regression coefficients, which are maximum likelihood estimates derived from maximizing the likelihood of the observed data. This same principle applies even when there are missing values in the data, as the estimation can still be based on the likelihood of the observed values. However, when missing data is present, the likelihood of the observed data becomes more complex compared to the typical data analysis scenarios, and finding an estimate that maximizes this likelihood doesn’t have a straightforward solution. To tackle the challenge, Dempster et al. (1977) proposed the Expectation-Maximization (EM) algorithm for incomplete data. The main idea and algorithm are explained in the following context.
@@
The two steps are repeated until some convergence conditions are met. Finally, the missing values are filled by using the estimated $\theta$. See the example of multivariate Gaussian in Section 11.6 of (Murphy, 2012).
@@ 5.5.1 Low-Rank Matrix Completion
Low-rank matrix completion (LRMC) (Srebro and Shraibman, 2005; Candès and Recht, 2009; Wen et al., 2012; Hardt, 2014) addresses the problem of recovering missing entries in a matrix under the assumption that the true matrix is low-rank. Formally, a matrix is low-rank if its rank (the dimension of the vector space spanned by its rows or columns) is significantly smaller than its dimensions. An equivalent characterization is that the matrix has a rapidly decaying singular value spectrum, with only a few non-zero or significant singular values. Consider a matrix $\mathbf{X}\in\mathbb{R}^{m\times n}$, where $m<n$, the singular value decomposition (SVD) is denoted as $\mathbf{X}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$, where $\mathbf{U}\in\mathbb{R}^{m\times m}$ and $\mathbf{V}\in\mathbb{R}^{n\times n}$ are orthonormal matrices, i.e., $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{m}$ and $\mathbf{V}^{\top}\mathbf{V}=\mathbf{I}_{n}$, $\bm{\Sigma}=(\text{diag}(\sigma_{1},\sigma_{2},\ldots,\sigma_{m}),\bm{0})$, and $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{m}$ are singular values. The rank of $\mathbf{X}$ is defined as $\text{rank}(\mathbf{X}):=\sum_{i=1}^{m}\mathbbm{1}(\sigma_{i}\neq 0)$. It has been found that many data matrices in real applications are low-rank ($\text{rank}(\mathbf{X})\ll m$) or can be well approximated by a low-rank matrix, i.e., many singular values are close to zero. Figure 10(a) shows a low-rank matrix of size $5\times 5$, where the rank is 2. Figure 10(b) shows the corresponding LRMC task.
(a) A rank-2 low-rank matrix of size $5\times 5$
(b) Completion of the low-rank matrix
Figure 10: A toy example of LRMC.
@@
nuclear norm $\|\mathbf{X}\|_{\ast}=\sum_{i=1}^{m}\sigma_{i}(\mathbf{X})$ (Candès and Recht, 2009)
truncated nuclear norm $\|\mathbf{X}\|_{\text{TN}}=\sum_{i=r+1}^{m}\sigma_{i}(\mathbf{X})$ (Hu et al., 2012)
weighted nuclear norm $\|\mathbf{X}\|_{\text{WN}}=\sum_{i=1}^{m}w_{i}\sigma_{i}(\mathbf{X})$ (Gu et al., 2014)
Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{Sp}}=(\sum_{i=1}^{m}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (Nie et al., 2012)
weighted Schatten-$p$ quasi-norm $\|\mathbf{X}\|_{\text{WSp}}=(\sum_{i=1}^{m}w_{i}\sigma_{i}^{p}(\mathbf{X}))^{1/p}$ (Xie et al., 2016b)
generalized non-convex penalty $R(\mathbf{X})=\sum_{i=1}^{m}g_{\lambda}(\sigma_{i}(\mathbf{X}))$ (Lu et al., 2014)
@@
factored nuclear norm $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{F}^{2}+\|\mathbf{B}\|_{F}^{2})$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\ast}$ (Srebro et al., 2005)
max norm $R(\mathbf{A},\mathbf{B})=(\max_{i}\left\|\mathbf{A}_{i:}\right\|_{2})(\max_{j}\left\|\mathbf{B}_{:j}\right\|_{2})$ Srebro and Shraibman (2005)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{*}+\|\mathbf{B}\|_{*})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (Shang et al., 2016)
factored Schatten-$1/2$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2}(\|\mathbf{A}\|_{2,1}+\left\|\mathbf{B}^{T}\right\|_{2,1})$ $\min R^{2}(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{\mathrm{S}_{1/2}}$ (Fan et al., 2019)
factored Schatten-$2q/(2+q)$ $R(\mathbf{A},\mathbf{B})=\frac{1}{q}\|\mathbf{A}\|_{2,q}^{q}+\frac{\alpha}{2}\left\|\mathbf{B}^{T}\right\|_{F}^{2}$ $\min R(\mathbf{A},\mathbf{B})=c\|\mathbf{X}\|_{S_{2q/(2+q)}}^{2q/(2+q)}$ (Fan et al., 2019)
factored Schatten-$p$ $R(\mathbf{A},\mathbf{B})=\frac{1}{2^{p}}\sum_{i=1}^{d}\left(\left\|\mathbf{A}_{:i}\right\|_{2}^{2}+\left\|\mathbf{B}_{i:}\right\|_{2}^{2}\right)^{p}$ $\min R(\mathbf{A},\mathbf{B})=\|\mathbf{X}\|_{S_{p}}^{p}$ (Giampouras et al., 2020)
@@
Beyond general low-rank matrices, several LRMC methods are designed for specific types. Notable examples include distance matrix completion (Alfakih et al., 1999), covariance matrix completion (Hippert-Ferrer et al., 2022), kernel matrix completion (Tsuda et al., 2003), and one-bit matrix completion (Davenport et al., 2014). In the latter case, the observed data are binary (1-bit) measurements rather than precise numerical values. This approach is valuable in scenarios with highly quantized data or binary feedback, such as in recommender systems and social networks.
@@ 5.5.2 High-Rank Matrix Completion
where $f^{(1)},\ldots,f^{(s)}$ are different nonlinear functions. This forms a matrix $\mathbf{X}=[\mathbf{X}^{(1)},\mathbf{X}^{(2)},\ldots,\mathbf{X}^{(s)}]$ of size $m\times n$, where $\mathbf{X}^{(j)}=[\mathbf{x}_{1}^{(j)},\ldots,\mathbf{x}^{(j)}_{n_{j}}]\in\mathbb{R}^{m\times n_{j}}$ and $n=\sum_{j=1}^{s}n_{j}$. It can be seen that (14) degrades to (13) if $s=1$ and degrades to (12) if all $f^{(j)}$ are linear. Figure 11 provides synthetic examples for (12),(13), and (14), respectively. The completion of the matrix produced by (14) poses the greatest challenge. Figure 12 shows an example of completing a matrix generated by (13).
Figure 11: Examples (Fan et al., 2020a) of data forming high-rank matrices in 3D space (left: union of subspaces; middle: one nonlinear manifold; right: union of nonlinear manifolds).
Figure 12: An example of HRMC on synthetic data. It shows the performance of the KFMC algorithm of (Fan and Udell, 2019b) in different iterations. The red points denote the true complete data, while the blue points denote the recovered data.
In recent years, a few high-rank matrix completion (HRMC) methods have been proposed (Eriksson et al., 2011; Alameda-Pineda et al., 2016; Elhamifar, 2016; Fan and Cheng, 2018; Fan et al., 2020a; Le Morvan et al., 2020). They can be organized into three categories: subspace-based methods, kernel-based methods, and deep learning-based methods.
@@
These methods were proposed for the data given by (12) and usually learn the subspaces from the incomplete data (Eriksson et al., 2011; Fan et al., 2020b) or exploit the self-expressive model (Yang et al., 2015; Elhamifar, 2016; Fan and Chow, 2017b, a), where each data point can be well represented by other data points drawn from the same subspace. The self-expressive model (Elhamifar and Vidal, 2013) can be formulated as
@@
where $\mathbf{S}\in\mathbb{R}^{n\times n}$ is the coefficient matrix and is usually assumed to be sparse. Given an incomplete data matrix $\tilde{\mathbf{X}}$, Fan and Chow (2017a) proposed the following matrix completion method:
@@
These methods (Alameda-Pineda et al., 2016; Ongie et al., 2017; Fan and Chow, 2018; Fan and Udell, 2019b; Fan et al., 2020a; Ongie et al., 2021) take advantage of kernel techniques that can convert a nonlinear problem in the original data space to a linear problem in the high-dimensional feature space. Ongie et al. (2017) and Fan and Chow (2018) proposed to solve
@@
Fan and Udell (2019b) showed that $\text{rank}(\phi(\mathbf{X}))=\min\{\binom{m+q}{q},n,s\binom{d+pq}{pq}\}$ if $f^{(j)}$ are $p$-order polynomials and $\phi$ is the feature map of a $q$-order polynomial kernel. When $n$ is sufficiently large and $p$ and $q$ are relatively small, the rank of $\phi(\mathbf{X})$ is low compared to its dimensions. Fan and Udell (2019b) introduced a kernel factorization method for HRMC and an online algorithm, which are more efficient than the methods proposed in (Ongie et al., 2017; Fan and Chow, 2018).
@@
These methods (Fan and Chow, 2017c; Fan and Cheng, 2018; Seyedi et al., 2023) are based on neural networks and will be detailed in Section 5.6.